OpenStudy (anonymous):

What is the vertex of y = |x + 1| ahhh i know its easy, but i'm getting it wrong.. :(

jimthompson5910 (jim_thompson5910):

y = |x + 1| y = 1*|x - (-1)| + 0 Now compare this to the form y = a*|x-h| + k

jimthompson5910 (jim_thompson5910):

what is h? what is k?

OpenStudy (anonymous):

-1 , 0

jimthompson5910 (jim_thompson5910):

good, so the vertex is (-1,0)

jimthompson5910 (jim_thompson5910):

since the vertex in general is (h,k)

OpenStudy (anonymous):

thanks:) i get it now @jim_thompson5910

jimthompson5910 (jim_thompson5910):

that's great

OpenStudy (anonymous):

you can also get it by solving y=x+1 and y=-x-1

OpenStudy (anonymous):

oh okay, yeah thanks:)